The Treasury bill rate is 4%, and the expected return on the market portfolio is 14%. According to the capital asset pricing model:

a. What is the risk premium on the market?

b. What is the required return on an investment with a beta of 1.4? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)

c. If an investment with a beta of 0.7 offers an expected return of 9.0%, does it have a positive or negative NPV?

d. If the market expects a return of 12.0% from stock X, what is its beta? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

If the market expects a return of 12.0% from stock X, what is its beta? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Sol: (a) Risk Premium an market = Rm- Rf - 14% -4% b) Reguüsesl Retumon. Investment = R1 + (Rm Rpx Bola = 47+(14-4)%* 1.4 = 1